OpenStudy (anonymous):

Which value is not a zero of P(x) = x^3 + 3x^2 - x - 3? f 1 g -3 h -1 j 3

OpenStudy (solomonzelman):

After plugging in which value for x, will the p(x) NOT be equal to zero ?

OpenStudy (solomonzelman):

\[P(x) = x^3 + 3x^2 - x - 3\]\[P(x) = x^2(x + 3) -( x + 3)\]\[P(x) = (x^2-1)(x + 3) \]this is what we get when we factor them. \[x^2-1=0~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~x+3=0\]\[(x-1)(x+1)=0~~~~~~~~~~~~~~~~~~x=-3\]\[x=±~1\]

OpenStudy (solomonzelman):

Which of the options is not zero ?

OpenStudy (anonymous):

3

OpenStudy (solomonzelman):

Yes, 3 is the only correct option :)

OpenStudy (anonymous):

thank you!

OpenStudy (solomonzelman):

Anytime !